OpenStudy (anonymous):

First clear the system of fractions then solve by substitution x/2-y/3=6/5 x/5-y/4=-13/10 HELP!

OpenStudy (praxer):

I will do the first, you deal with the second. :) x/2-y/3=6/5 =>(3x-2y)/6=6/5 now, cross multiply, =>15x-10y=36.